OpenStudy (anonymous):

A poorly tuned VW can accelerate from rest to speed of 28 m/s in 20 seconds. What is the average acceleration?

OpenStudy (michele_laino):

please, by definition, you have to apply this formula, in order to find the average acceleration a: \[a=\frac{ \Delta v }{ \Delta t }\]

OpenStudy (michele_laino):

where Delta_v is equals to the speed variation, and Delta_t is the time interval i which Delta_v is ooccurred

OpenStudy (michele_laino):

please try to substitute your numerical data into the formula above!

OpenStudy (godlovesme):

U still need help @hl147 :)

OpenStudy (godlovesme):

K so it started from rest which means 0m/s and reach at speed of 28m/20s Av= final speed/ initial speed/ final time-initial time Av= 28-0/ 20-0 Av= 28/20 Av= 1.4m/s Make sense? :)

OpenStudy (michele_laino):

@Godlovesme that's right! even if the unit of measure is m/sec^2, not m/sec

OpenStudy (godlovesme):

Yea my fault got it @hl147

OpenStudy (anonymous):

yes thank you! @Godlovesme

OpenStudy (godlovesme):

Anytime :)

OpenStudy (godlovesme):

Thanks for the medal! :)

OpenStudy (anonymous):

Your welcome!!